# A Survey on Large Language Model Acceleration based on KV Cache Management

### v2

@@ 5.1.2 Cross-layer Sharing
DHA (Decoupled Head Attention) [193] addresses redundancy in MHA and adaptively configures shared groups for key and value heads across layers, reducing KV cache requirements. Observing that clustering and fusing similar heads can reduce KV cache size without significant performance reduction, DHA designs a search, fusion, and continued pre-training framework that can progressively transform MHA checkpoints into DHA models through linear fusion of head parameters, preserving the pre-trained knowledge with small pre-training budget.
5.1.3 Summary and Future Directions
This section highlights innovative strategies for optimizing memory and computational efficiency through intra-layer grouping and cross-layer sharing mechanisms. However, several avenues for improvement remain. First, maintaining performance while optimizing efficiency, especially for precision-sensitive tasks, requires further investigation. Methods that implement radical grouping and sharing mechanisms may compromise the model fidelity for tasks requiring high precision. Second, scalability across diverse model architectures and sizes is essential. Works such as DHA [193] and LISA [190], which rely on specific architectural assumptions, may struggle to generalize to emerging LLMs or non-standard configurations. Third, the dynamics of attention across both time and layers are largely under-explored. Most existing methods rely on static or pre-determined grouping and sharing strategies, neglecting the temporal and contextual variations in attention patterns.
To address these challenges and unlock the full potential of attention optimization, future research should focus on the following aspects. First, developing universal frameworks for attention grouping and sharing that require minimal retraining to enhance adaptability and usability. Second, synergistic integration with other optimization techniques, such as quantization and pruning, has significant potential to achieve even greater efficiency gains. While some works like CLLA [192] have begun to address these opportunities, more exploration could be carried out to unlock new levels of efficiency. Third, more dynamic and temporal modeling could be leveraged to adaptively adjust grouping and sharing during runtime to better capture the contextual requirements of different tasks and sequences. Finally, a deeper understanding of the downstream impacts of these techniques on fine-tuning and transfer learning is crucial for their effective application in real-world scenarios.
@@ 5.2 Architecture Alteration
This section explores architectural modifications to optimize KV cache usage. We categorize these methods into two subsections: methods that refine the attention mechanism for KV cache efficiency (Sec. 5.2.1), and methods that introduce structural changes for better KV management (5.2.2). Many of these works build upon the broader landscape of efficient attention mechanisms (e.g., Linear Transformer [202], Performer [203], LinFormer [204], etc.). Since our focus lies on methods directly impacting KV cache handling, for a comprehensive overview of efficient attention mechanisms, we refer readers to dedicated surveys [45]. The summary of architecture alteration for KV reuse is listed in Tab. VIII.
@@ 5.2.1 Enhanced Attention
DeepSeek-V2 [27] introduced Multi-Head Latent Attention (MLA) that adopts a low-rank KV joint compression mechanism, replacing the full KV cache with compressed latent vectors. The model adopts trainable projection and expansion matrices to do the compression. This compression mechanism significantly reduces the memory requirement of the KV cache and allows the model to handle sequences up to 128K tokens.
FLASH [184] incorporates the Gated Attention Unit (GAU) to replace the MHA mechanism in traditional transformers. GAU utilizes a single-head attention mechanism with gating functions that selectively modulates importance in information flow. FLASH employs a linear approximation method for attention computation through GAU module, which makes the model efficiently handle long contexts without the quadratic scaling of traditional self-attention, thus mitigating heavy KV cache issues.
@@ 5.2.2 Augmented Architecture
YOCO [180] builds a decoder-decoder architecture composed of two modules: a self-decoder and a cross-decoder. The self-decoder efficiently encodes global key-value caches, while the cross-decoder reuses these caches via cross-attention. This design ensures that key-value pairs are only cached once, substantially reducing GPU memory usage while maintaining global attention capabilities. YOCO’s computation flow also enables the prefilling to early exit, allowing faster prefill stages without altering the final output.
CEPE [181] interleaves additional cross-attention layers between the self-attention and feed-forward layers in the decoder model. It employs a small encoder to process long inputs chunk-by-chunk to encoded representations as cross-attention layers’ inputs. In this way, CEPE can prevent the needs for KV cache for every token and reduce computational cost by processing contexts in parallel. This also facilitates an existing LLMs to expand its contexts while preserving the scalability and generalizability.
XC-Cache [182] also utilizes an encoder to interleave cross-attention layers within existing self-attention layers in pre-trained decoder-only models to prevent explicit prompt caching. The encoder processes the context and converts it into a compact set of key-value pairs that summarize the essential information. It also finds that pre-trained causal decoders can be used to replace an encoder for representations extraction, further reducing the training costs on additional encoder.
@@ 5.3 Non-Transformer Architecture
While transformers are struggling with KV cache issues, researchers have revisited principles from traditional sequential architectures, such as recurrent neural networks (RNNs) [205], which inherently process sequences without the need for explicit KV caches. Inspired by the lightweight and memory-efficient design of RNNs and efficient attention mechanisms, non-transformer architectures [206, 207, 208, 209, 177, 176] have emerged, such as Mamba [177] and RWKV [176], offering promising alternatives. While there are a large type of new architectures, we only list methods associated with KV optimization. For further understanding to efficient non-transformer works, please refer to these surveys [45, 210, 211, 212]. The summary of non-transformer is listed in Tab. IX.
@@ 5.3.1 Adaptive Sequence Processing Architectures
RetNet [178] introduces Retentive Network that combines elements of recurrence and attention, presenting a novel retention mechanism for sequence modeling that offers training parallelism, low-cost inference, and scalable performance together. The proposed Multi-scale Retention Module (MSR) enables support to multiple computation paradigms: the parallel representation is similar to self-attention that adds support to casual masks and parallel training. The recurrent representation is similar to RNN that allows low-cost inference by maintaining state across sequence decoding. The chunkwise recurrent representation constructs a hybrid form to the former representations to further enables handling long sequences. These combined characteristics position RetNet as a strong alternative to transformers without a heavy KV cache mechanism.
MCSD [179] features the new block called Multi-Channel Slope and Decay, which is made up of two sections: The slope section can capture local features across short temporal spans, and the decay section can capture global features across long temporal spans. The sections are fused through element-wise operations. During inference, the process would be reformat into a recurrent representation, allowing both spatial and temporal efficiency, minimizing the need for maintaining a large KV cache.
@@ 5.3.2 Hybrid Architecture
With these non-transformer architecture, some methods construct mixed models to alleviate KV cache necessities while keeping some peculiarities and merits of the self-attention mechanism.
MixCon [173] introduces a new architecture called Conba. Inspired by control theory, the Conba layer incorperates a feedback and adaptive control mechanism that can adapt to different sequence-modeling tasks and requirements dynamically with good computational efficiency. Furthermore, MixCon integrates the Mixture of Experts (MoE) module, which dynamically selects the most relevant experts to process parts of the sequence. Combining the transformer layer, the Conba layer, and the MoE module, MixCon constructs a hybrid model with good balance between attention effectiveness and computational efficiency and significantly reduces the total size of the KV cache.
@@
RecurFormer [175] argues that not all transformer heads need to participate in the self-attention mechanism. The work recognizes that certain attention heads show recency-aware behavior which focus on local and short-range dependencies, dissipate the computation resource but gives little contribution. After identifying these heads, RecurFormer replaces them with the Mamba components, achieving straightforward KV cache reduction.
@@ 5.3.3 Summary and Future Directions
By exploring non-transformer modules such as recurrent and hybrid designs, these methods have introduced novel paradigms that balance performance with computational efficiency, and also alleviate the KV cache issues in traditional transformer architectures. Future research should focus on several key areas. First, improving the scalability of recurrent architectures, such as RWKV [176] and Mamba [177], remains critical. Although these methods reduce memory and computational costs, their performance in capturing ultra-long-range dependencies lags behind transformers. Second, hybrid designs such as MixCon [173] and GoldFinch [174] highlight the potential of integrating diverse modules, yet their complexity introduces challenges in training stability and interpretability. Third, the overall generalization capabilities and robustness of non-transformer architectures, while efficient, need require further exploration for diverse input modalities.
@@ 6 System-level Optimization
Design (Sec. 6.3.3), snode [NEO [216], FastDecode [217], FlexInfer [218], InfiniGen [139], Pensieve [219], FastServe [220], PartKVRec [221], wnode_1] ] [I/O-based
Design (Sec. 6.3.2), snode [FlashAttention [145], Bifurcated Attention [222], PartKVRec [221], HCache [223], Cake [224], FastSwitch [225], wnode_1] ] [Single/Multi-GPU Design (Sec. 6.3.1), snode [HydraGen [226], DeFT [227], vLLM [144], ORCA [228], DistServe [229], Multi-Bin Batching [230], Tree Attention [231], wnode_1] ] ] [Scheduling (Sec. 6.2), snode [Layer-specific and Hierarchical Scheduling (Sec. 6.2.3), snode [LayerKV [232], CachedAttention [233], ALISA [234], LAMPS [235], wnode_1] ] [Preemptive and Fairness-oriented Scheduling (Sec. 6.2.2), snode [FastServe [220], FastSwitch [225], wnode_1] ] [Prefix-aware Scheduling (Sec. 6.2.1), snode [BatchLLM [236], RadixAttention [237], wnode_1] ] ] [Memory
Management (Sec. 6.1), snode [Prefix-aware Design (Sec. 6.1.2), snode [ChunkAttention [238], MemServe [239], wnode_1] ] [Architectural Design (Sec. 6.1.1), snode [vLLM [144], vTensor [218], LeanKV [112], wnode_1] ] ] ]
@@
Recent system-level optimizations for KV cache in LLM inference can be broadly categorized into three main directions: memory management (Sec. 6.1), scheduling strategies (Sec. 6.2), and hardware-aware designs (Sec. 6.3). These complementary approaches collectively demonstrate the rich design space for system-level optimizations in LLM inference, each addressing different aspects of the performance, efficiency, and resource utilization challenges. The Taxonomy of the system-level optimization is in Fig. 6.
@@ 6.1 Memory Management
Recent advances in KV cache memory management for large language model (LLM) inference reveal three distinct approaches aimed at enhancing memory efficiency. Architectural designs, exemplified by vLLM with PagedAttention [144] and vTensor [218], adapt classical operating system principles to create flexible, dynamic memory allocation systems that optimize the use of physical memory through sophisticated mapping and virtual memory abstractions. Prefix-aware designs like ChunkAttention [238] and MemServe [239] further refine this approach by organizing data structures to enable efficient cache deduplication and sharing of common prefixes, thereby improving both memory utilization and computational efficiency. Together, these innovations illustrate the potential for significant enhancements in LLM serving via memory management.
@@ 6.1.1 Architectural Design
This approach is further enhanced by vTensor [218], which introduces a virtual memory abstraction that decouples computation from defragmentation through three key components: the vTensor Scheduler which generates memory management policies based on meta information, the vTensor Operation which translates these policies into CUDA VMM operations, and the vTensor Pool which maintains virtual tensor mappings. VTS processes instructions and creates policies based on memory state tracking, while VTO executes these policies through asynchronous GPU operations. VTP completes the cycle by managing virtual tensor storage and updating meta information for subsequent memory operations.
LeanKV [112] combines unified paging with heterogeneous quantization and dynamic sparsity mechanisms. It implements Hetero-KV quantization to store keys and values at different precisions, complemented by a per-head dynamic sparsity mechanism that adapts memory allocation based on token importance across different attention heads and requests. To efficiently execute these strategies, LeanKV [112] introduces an advanced on-GPU memory management system featuring three key components: unified paging for flexible memory organization, a circular free page list for efficient coordination, and a bidirectional page table for minimal metadata overhead.
@@ 6.1.2 Prefix-aware Design
These approaches often complement each other, suggesting potential benefits in combining multiple strategies. For instance, LeanKV [112]’s integration of compression with page-based management and MemServe [239]’s combination of distributed memory management with prefix-aware caching demonstrate the effectiveness of hybrid approaches. The diversity of these solutions reflects both the complexity of KV cache management and the rich opportunity space for continued innovation in optimizing LLM inference systems. Tab.X provides a comparison of various memory management techniques for KV Cache, highlighting key features such as paged memory, virtual memory, dynamic sparsity, prefix sharing, and distributed memory.
6.1.3 Summary and Future Directions
Future work should advance memory management innovations through multiple synergistic directions: investigating adaptive memory hierarchies that dynamically adjust to workload patterns and resource constraints, exploring novel compression techniques that preserve quick access while reducing memory footprint, developing intelligent prefetching mechanisms that anticipate and preload frequently accessed cache entries, researching hardware-aware optimization strategies that leverage emerging memory technologies like computational storage and processing-in-memory units, and designing distributed cache coherence protocols that efficiently maintain consistency across multiple inference nodes. Additionally, the exploration of machine learning-based approaches could enable predictive memory allocation that learns from historical access patterns, while the investigation of specialized data structures could yield more efficient prefix detection and sharing mechanisms. These advancements, combined with research into heterogeneous memory systems that intelligently coordinate different memory types based on access patterns and performance requirements, would significantly enhance the scalability and efficiency of LLM inference systems across diverse deployment scenarios.
@@ 6.2.1 Prefix-aware Scheduling
RadixAttention [237] builds around a radix tree structure, replacing traditional FCFS scheduling with an intelligent cache-aware approach that prioritizes requests based on matched prefix lengths. It implements dynamic memory management where cached tokens and running requests share the same memory pool, controlled by an LRU eviction policy that strategically removes leaf nodes while preserving valuable ancestor prefixes. This is complemented by a reference counting mechanism that prevents eviction of actively used cache entries during continuous batching while enabling efficient memory reclamation when nodes become unused.

### v3

@@ 5.2 Architecture Alteration
This section explores architectural modifications to optimize KV cache usage. We categorize these methods into two subsections: methods that refine the attention mechanism for KV cache efficiency (Sec. 5.2.1), and methods that introduce structural changes for better KV management (Sec. 5.2.2). Many of these works build upon the broader landscape of efficient attention mechanisms (e.g., Linear Transformer [213], Performer [214], LinFormer [215], etc.). Since our focus lies on methods directly impacting KV cache handling, for a comprehensive overview of efficient attention mechanisms, we strongly refer readers to dedicated surveys [46]. The summary of architecture alteration for KV reuse is clearly listed in Tab. VIII.
@@ 5.2.1 Enhanced Attention
DeepSeek-V2 [28] introduced Multi-Head Latent Attention (MLA) that adopts a low-rank KV joint compression mechanism, replacing the full KV cache with compressed latent vectors. The model adopts trainable projection and expansion matrices to do the compression. This compression mechanism significantly reduces the memory requirements of the KV cache and enables the model to handle sequences of up to 128K tokens.
FLASH [195] incorporates the Gated Attention Unit (GAU) to replace the MHA mechanism in traditional transformers. GAU employs a single-head attention mechanism with gating functions that selectively modulate the importance of information flow. FLASH employs a linear approximation method for attention computation through GAU module, which makes the model efficiently handle long contexts without the quadratic scaling of traditional self-attention, thus mitigating heavy KV cache issues.
@@ 5.2.2 Augmented Architecture
YOCO [191] builds a novel decoder-decoder architecture composed of two modules: a self-decoder and a cross-decoder. The self-decoder efficiently encodes global key-value caches, while the cross-decoder reuses these caches via cross-attention. This design ensures that key-value pairs are only cached once, substantially reducing GPU memory usage while maintaining global attention capabilities. Furthermore, YOCO’s computation flow also enables the prefilling to early exit, allowing faster prefill stages without altering the final output.
CEPE [192] interleaves additional cross-attention layers between the self-attention and feed-forward layers in the decoder model. It employs a small encoder to process long inputs chunk-by-chunk to encoded representations as cross-attention layers’ inputs. In this way, CEPE can prevent the need for KV cache for every token and reduce computational cost by processing contexts in parallel. This also facilitates an existing LLM to expand its contexts while preserving the scalability and generalizability.
XC-Cache [193] also utilizes an encoder to interleave cross-attention layers within existing self-attention layers in pre-trained decoder-only models to prevent explicit prompt caching. The encoder processes the context and converts it into a compact set of key-value pairs that summarize the essential information. It also finds that pre-trained causal decoders can be used to replace an encoder for the representations extraction, further reducing the training costs on an additional encoder.
@@ 5.3 Non-Transformer Architecture
While transformers are struggling with KV cache issues, researchers have revisited principles from traditional sequential architectures, such as recurrent neural networks (RNNs) [216], which inherently process sequences without the need for explicit KV caches. Inspired by the lightweight and memory-efficient design of RNNs and efficient attention mechanisms, non-transformer architectures [217, 218, 219, 220, 188, 187] have emerged, such as Mamba [188] and RWKV [187], offering promising alternatives. While there are a large type of new architectures, we only list methods associated with KV optimization. For further understanding of efficient non-transformer works, please refer to these surveys [46, 221, 222, 223]. The summary of non-transformer is listed in Tab. IX.
@@ 5.3.1 Adaptive Sequence Processing Architectures
RetNet [189] introduces the Retentive Network, which combines elements of recurrence and attention, presenting a novel retention mechanism for sequence modeling that enables training parallelism, low-cost inference, and scalable performance. The proposed Multi-scale Retention Module (MSR) supports multiple computation paradigms: the parallel representation, similar to self-attention, allows for causal masking and parallel training; the recurrent representation, akin to RNNs, enables low-cost inference by maintaining state across sequence decoding; and the chunkwise recurrent representation constructs a hybrid of the previous two approaches, further enhancing the ability to effectively handle long sequences.
MCSD [190] introduces a new block called Multi-Channel Slope and Decay, which consists of two sections: the slope section, which captures local features across short temporal spans, and the decay section, which captures global features across long temporal spans. These sections are fused through element-wise operations. During inference, the process is reformulated into a recurrent representation, allowing for both spatial and temporal efficiency and minimizing the need to maintain a large KV cache.
Figure 9: Computer server architecture.
@@ 5.3.2 Hybrid Architecture
With these non-transformer architectures, some methods construct mixed models to alleviate KV cache necessities while keeping some peculiarities and merits of the self-attention mechanism.
MixCon [184] introduces a new architecture called Conba. Inspired by control theory, the Conba layer incorporates a feedback and adaptive control mechanism that can adapt to different sequence-modeling tasks and requirements dynamically with good computational efficiency. Furthermore, MixCon integrates the Mixture of Experts (MoE) module, which dynamically selects the most relevant experts to process parts of the sequence. Combining the transformer layer, the Conba layer, and the MoE module, MixCon constructs a hybrid model with a good balance between attention effectiveness and computational efficiency and significantly reduces the total size of the KV cache.
@@
RecurFormer [186] argues that not all transformer heads need to participate in the self-attention mechanism. The authors observe that certain attention heads exhibit recency-aware behavior, focusing on local and short-range dependencies. These heads consume computational resources but contribute little to overall performance. After identifying such heads, RecurFormer replaces them with Mamba components, resulting in straightforward KV cache reduction.
@@ 5.3.3 Summary and Future Directions
By exploring non-transformer modules such as recurrent and hybrid designs, these methods have introduced novel paradigms that balance performance with computational efficiency, and also alleviate the KV cache issues in traditional transformer architectures. Future research should focus on several key areas. First, improving the scalability of recurrent architectures, such as RWKV [187] and Mamba [188], remains critical. Although these methods reduce memory and computational costs, their performance in capturing ultra-long-range dependencies lags behind transformers. Second, hybrid designs such as MixCon [184] and GoldFinch [185] highlight the potential of integrating diverse modules, yet their complexity introduces challenges in training stability and interpretability. Third, the overall generalization capabilities and robustness of non-transformer architectures need exploration for diverse input modalities.
@@ 6 System-level Optimization
Design (Sec. 6.3.3), snode [NEO [227], FastDecode [228], FlexInfer [229], InfiniGen [144], Pensieve [230], FastServe [231], PartKVRec [232], HeadInfer [233],APEX [234], wnode_1] ] [I/O-based
Design (Sec. 6.3.2), snode [FlashAttention [150] [151] [152], Bifurcated Attention [235], PartKVRec [232], HCache [236], Cake [237], FastSwitch [238], wnode_1] ] [Single/Multi-GPU Design (Sec. 6.3.1), snode [HydraGen [239], DeFT [240], vLLM [149], ORCA [241], DistServe [242], Multi-Bin Batching [243], Tree Attention [244],gLLM [245],FairKV [246], MELL [247], wnode_1] ] ] [Scheduling (Sec. 6.2), snode [Layer-specific and Hierarchical Scheduling (Sec. 6.2.3), snode [LayerKV [248], CachedAttention [249], ALISA [250], LAMPS [251], Apt-Serve [252],FGOS [253], wnode_1] ] [Preemptive and Fairness-oriented Scheduling (Sec. 6.2.2), snode [FastServe [231], FastSwitch [238], FlowKV [254], wnode_1] ] [Prefix-aware Scheduling (Sec. 6.2.1), snode [BatchLLM [255], RadixAttention [256], Echo [257], wnode_1] ] ] [Memory
Management (Sec. 6.1), snode [Prefix-aware Design (Sec. 6.1.2), snode [ChunkAttention [258], MemServe [259], FlashForge [260], wnode_1] ] [Architectural Design (Sec. 6.1.1), snode [vLLM [149], vTensor [229], LeanKV [116], eLLM [261], Apt-Serve [252], wnode_1] ] ] ]
@@
As shown in Fig. 9, existing computing server systems consist of various components, such as GPUs, CPUs, and memory storage. Optimizing LLM acceleration via KV cache at the system level is both practical and intriguing, considering the different communication and data exchange mechanisms. Recent system-level optimizations for KV cache in LLM inference can be broadly categorized into three main directions: memory management (Sec. 6.1), scheduling strategies (Sec. 6.2), and hardware-aware designs (Sec. 6.3). These complementary approaches collectively demonstrate the rich design space for system-level optimizations in LLM inference, each addressing different aspects of the performance, efficiency, and resource utilization challenges. The taxonomy of the system-level optimization is in Fig. 10.
@@ 6.1 Memory Management
Recent advances in KV cache memory management for large language model (LLM) inference reveal three distinct approaches aimed at enhancing memory efficiency. Architectural designs, exemplified by vLLM with PagedAttention [149] and vTensor [229], adapt classical operating system principles to create flexible, dynamic memory allocation systems that optimize the use of physical memory through sophisticated mapping and virtual memory abstractions. Prefix-aware designs like ChunkAttention [258] and MemServe [259] further refine this approach by organizing data structures to enable efficient cache de-duplication and sharing of common prefixes, thereby improving both memory utilization and computational efficiency. Together, these innovations illustrate the potential for significant enhancements in LLM serving via memory management.
@@ 6.1.1 Architectural Design
This approach is further enhanced by vTensor [229], which introduces a virtual memory abstraction that decouples computation from defragmentation through three key components: the vTensor Scheduler, which generates memory management policies based on meta information, the vTensor Operation, which translates these policies into CUDA VMM operations, and the vTensor Pool, which maintains virtual tensor mappings. VTS processes instructions and creates policies based on memory state tracking, while VTO executes these policies through asynchronous GPU operations. VTP completes the cycle by managing virtual tensor storage and updating meta information for subsequent memory operations.
LeanKV [116] combines unified paging with heterogeneous quantization and dynamic sparsity mechanisms. It implements Hetero-KV quantization to store keys and values at different precisions, complemented by a per-head dynamic sparsity mechanism that adapts memory allocation based on token importance across different attention heads and requests. To efficiently execute these strategies, LeanKV [116] introduces an advanced on-GPU memory management system featuring three key components: unified paging for flexible memory organization, a circular free page list for efficient coordination, and a bidirectional page table for minimal metadata overhead. DMS (Dynamic Memory Sparsification) [262] is a method for sparsifying KV caches while maintaining better accuracy than training-free sparse attention.
eLLM [261] is a memory management framework that draws inspiration from the classical memory ballooning technique used in operating systems. The key components of eLLM include: a). Virtual Tensor Abstraction: This layer decouples the virtual address space of tensors from the physical GPU memory, allowing for the creation of a unified and flexible memory pool; b). Elastic Memory Mechanism: This dynamic system adjusts memory allocation by inflating and deflating the memory usage at runtime. It leverages the CPU memory as an extensible buffer to facilitate this flexible memory management; c). Lightweight Scheduling Strategy: eLLM employs SLO-aware policies within this scheduling component to optimize memory utilization and effectively balance performance trade-offs while adhering to strict SLO constraints. Apt-Serve [252] features a new hybrid cache scheme that combines KV cache with a memory-efficient hidden cache for reusable input hidden state vectors, allowing large batch sizes and improving request concurrency.
@@ 6.1.2 Prefix-aware Design
FlashForge [260] delivers a novel shared-prefix attention kernel that optimizes memory hierarchy and exploits both intra-block and inter-block parallelism and a comprehensive workload balancing mechanism that efficiently estimates cost, divides tasks, and schedules execution.
@@
FlashForge [260]
These approaches often complement each other, suggesting potential benefits in effectively combining multiple strategies. For instance, LeanKV [116]’s integration of compression with page-based management and MemServe [259]’s combination of distributed memory management with prefix-aware caching demonstrates the effectiveness of hybrid approaches. The diversity of these solutions reflects both the complexity of KV cache management and the rich opportunity space for continued innovation in optimizing LLM inference systems. Tab.X provides a comparison of various memory management techniques for KV Cache, highlighting key features such as paged memory, virtual memory, dynamic sparsity, prefix sharing, and distributed memory.
6.1.3 Summary and Future Directions
Future memory management work should prioritize adaptive hierarchies, novel compression techniques, intelligent prefetching, and hardware-aware optimizations utilizing new memory technologies, alongside efficient distributed cache coherence. Exploring machine learning for predictive allocation, specialized data structures, and heterogeneous memory systems is also vital for enhancing LLM inference scalability and efficiency.
@@ 6.2.1 Prefix-aware Scheduling
RadixAttention [256] builds around a radix tree structure, replacing traditional FCFS scheduling with an intelligent cache-aware approach that prioritizes requests based on matched prefix lengths. It implements dynamic memory management where cached tokens and running requests share the same memory pool, controlled by an LRU eviction policy that strategically removes leaf nodes while preserving valuable ancestor prefixes. This is complemented by a reference counting mechanism that prevents eviction of actively used cache entries during continuous batching while enabling efficient memory reclamation when nodes eventually become unused.
Echo [257] is a collaborative online-offline task serving system that includes a scheduler and a KV cache manager. The scheduler and KV cache manager work tightly to maximize the throughput of offline tasks, while the estimator further predicts execution time to ensure online task SLOs. The scheduler leverages the batch information of last iteration to reduce the search space for finding the optimal schedule. The KV cache manager sets the priority of the KV cache based on the type of tasks and the opportunity of prefix sharing to reduce the recomputation.